### v3

# PromptGuard: Soft Prompt-Guided Unsafe Content Moderation for Text-to-Image Models

Abstract
Recent text-to-image (T2I) models have exhibited remarkable performance in generating high-quality images from text descriptions. However, these models are vulnerable to misuse, particularly generating not-safe-for-work (NSFW) content, such as sexually explicit, violent, political, and disturbing images, raising serious ethical concerns. In this work, we present PromptGuard, a novel content moderation technique that draws inspiration from the system prompt mechanism in large language models (LLMs) for safety alignment. Unlike LLMs, T2I models lack a direct interface for enforcing behavioral guidelines. Our key idea is to optimize a safety soft prompt that functions as an implicit system prompt within the T2I model’s textual embedding space. This universal soft prompt ($P^{*}$) directly moderates NSFW inputs, enabling safe yet realistic image generation without altering the inference efficiency or requiring proxy models. We further enhance its reliability and helpfulness through a divide-and-conquer strategy, which optimizes category-specific soft prompts and combines them into holistic safety guidance. Extensive experiments across five datasets demonstrate that PromptGuard effectively mitigates NSFW content generation while preserving high-quality benign outputs. PromptGuard achieves 3.8 times faster than prior content moderation methods, surpassing eight state-of-the-art defenses with an optimal unsafe ratio down to 5.84%.
I Introduction
Text-to-image (T2I) models, like Stable Diffusion [1], enable realistic image generation from text prompts. However, their misuse for generating not-safe-for-work (NSFW) content (e.g., sexual and violent images) raises significant ethical concerns [2, 3, 4], including the spread of harmful content like AI-generated child sexual abuse material [5] and politically manipulative imagery [6]. Effective defense mechanisms for T2I services are urgently needed.
Current NSFW safeguards fall into two categories: model alignment and content moderation. Model alignment (e.g., fine-tuning) directly modifies the T2I model to remove NSFW capabilities [7, 8, 9, 10, 11, 12], but can degrade performance on benign inputs [10, 13]. Content moderation uses external models to filter unsafe textual inputs [14] or visual outputs [15], or employs prompt modification using LLMs [16] to promote safer generation. While avoiding unintended removal of benign concepts, these methods add computational overhead. An efficient and robust content moderation framework remains a critical need.
Figure 1: Unlike existing moderation frameworks that rely on additional models to check or detoxify NSFW content, PromptGuard presents an efficient, universal soft prompt, $P_{*}$, inspired by the system prompt mechanism in LLMs, to directly moderates NSFW inputs and generate safe yet realistic content.
In this paper, we present PromptGuard, a novel T2I moderation technique that optimizes a soft prompt that works as a system prompt for safety to neutralize malicious contents in input prompts in an input-agnostic manner without affecting benign image generation quality and performance. As shown in Figure 1, our basic idea draws inspiration from the “system prompt” mechanism in LLMs, which has exhibited remarkable effectiveness in aligning output content with safe and ethical guidelines [17, 18] and our approach seeks to apply similar guidance in T2I settings.
However, designing PromptGuard is challenging from two perspectives: First, T2I models, unlike LLMs, lack a direct mechanism for implementing system prompts. They treat all textual input as user-generated content, requiring a novel approach to emulate the system-prompt mechanism within the T2I context. Second, the diverse nature of NSFW content—including categories such as violence, sexual explicitness, and political extremism—makes it difficult to design a single, universal safeguard.
To address the first challenge, we introduce a safety pseudo-word, optimized within the continuous embedding space of the T2I model’s text encoder. This soft prompt effectively steers both benign and NSFW prompts (e.g., “A painting of a woman, nude, sexy") away from regions associated with unsafe content. Moreover, we employ SDEdit[19] to transform unsafe images into safer counterparts, allowing PromptGuard to learn how to generate realistic, safe images from potentially harmful inputs. This approach contrasts with existing moderation methods[14, 15, 9] that often block or blur undesirable outputs. For the second challenge, we categorize NSFW content into four types: sexual, violent, political, and disturbing [20, 21]. Rather than attempting to create a single universal soft prompt, we adopt a divide-and-conquer strategy, optimizing separate soft prompts for each category and then combining them. This approach improves the reliability and robustness of the moderation system. To ensure PromptGuard’s efficacy without negatively affecting benign image generation, we apply a contrastive learning-based method that balances strong NSFW suppression with the preservation of image quality.
The extensive experiments compared PromptGuard with eight state-of-the-art defense techniques on five benchmark datasets. Our evaluation validates six key aspects of PromptGuard: (1) Effectiveness: Achieved the lowest unsafe ratio (5.84%) in a natural language setting, outperforming all baselines. (2) Universality: Ranked in the top two across all four NSFW categories. (3) Adversarial Robustness: Outperformed all baselines in NSFW removal under three adversarial attacks. (4) Efficiency: 3.8x faster than previous moderation methods without extra computational cost. (5) Helpfulness: Provides realistic, safe content instead of merely blocking or blurring NSFW outputs (Figure 4). (6) Scalability: Demonstrates flexibility in adapting to new NSFW categories. We also discuss limitations, future work, and have open-sourced our code on website [^1] Project Website: https://prompt-guard.github.io/ to foster further research in AI ethics.
Our contributions can be summarized as follows:
  • New Technique: We introduce the application of the system prompt concept to T2I models, using soft prompt optimization to achieve effective and lightweight content moderation.
  • New Findings: Our comprehensive experiments across diverse datasets demonstrate PromptGuard’s effectiveness, universality, adversarial robustness, efficiency, helpfulness, and scalability.
II Related Work
II-A Content Moderation
To ensure the safe use of T2I models, existing methods implement safety measures for both input and output. Latent Guard [22] filters input text by classifying embeddings, allowing only safe prompts to pass through. In contrast, Stable Diffusion V1.4’s default safety filter [15] detects and blocks any NSFW images at the output stage by blacking them out. POSI [16] fine-tunes a language model to rewrite unsafe prompts into safe alternatives before passing them to the diffusion model. Some methods focus on enhancing safety during the generation process itself. Safe Latent Diffusion [23] adjusts the diffusion process to steer the text-conditioned guidance vector away from unsafe areas in the embedding space. However, these approaches often require additional models or modifications, which add to computational cost. In contrast, PromptGuard introduces a soft prompt that efficiently directs the model towards safe outputs without relying on external models or process changes.
II-B Model Alignment
Another line of work directly fine-tunes models to enhance safety, rather than relying solely on additional guardrails. ESD [7] fine-tunes the diffusion model to direct the generative process away from undesired concepts, while UCE [8] modifies the text projection matrices to erase specific concepts from the model. Additionally, SafeGen [9] optimizes the self-attention layers to eliminate unsafe concepts in a text-agnostic manner. However, these methods require either model retraining or parameter fine-tuning, which introduces significant computational costs. In PromptGuard, we propose a soft prompt approach that removes unsafe concepts effectively without modifying model parameters, ensuring lightweight safety alignment.
III Background
III-A Text-to-Image (T2I) Generation
The success of denoising diffusion models, such as DDPM [24], has advanced text-to-image (T2I) models like Stable Diffusion (SD) and Latent Diffusion [25]. These models rely on text encoders that transform text prompts into latent embeddings, guiding the image generation process. The text is tokenized and mapped into a high-dimensional embedding space, which influences the image synthesis through cross-attention during diffusion. For instance, SD uses the CLIP text encoder, which improves upon the BERT encoder used in Latent Diffusion [26], benefiting from a larger training set (LAION-5B [27]) for more effective embeddings. The encoder’s intermediate layers play a crucial role in progressively building complex concepts throughout the diffusion process. Recent studies, like the Diffusion Lens [28], show that early layers capture basic objects, while deeper layers establish relationships between elements.
III-B System Prompt
A system prompt is a predefined instruction given to large language models (LLMs) to guide their behavior, tone, and responses, ensuring safety and mitigating risks such as bias or harmful outputs [29, 30]. By embedding ethical guidelines, system prompts prevent misleading responses without modifying model parameters [31]. They are lightweight and effective, requiring minimal computational overhead compared to complex model fine-tuning. Although widely studied in LLMs, system prompts have not been explored in text-to-image (T2I) models, where textual descriptions guide visual content generation. Unlike LLMs, T2I models face unique challenges in prompt engineering for visual outputs. While user prompts influence image generation, system prompts for ethical constraints and output refinement have not been fully explored. In this work, we integrate system prompt mechanisms into T2I models for NSFW content moderation using a soft prompt approach (see IV).
IV PromptGuard
IV-A Overview
Figure 2: Diagram of PromptGuard. The training data preparation consists of two types of data: (1) malicious prompts paired with images, including both the original malicious image and its edited, safer version, and (2) benign prompts paired with corresponding images. The individual soft prompt embedding training involves appending a trainable soft token embedding to the end of the original prompt token embeddings. Focusing on one unsafe category at a time, we train only the parameters of the soft token embedding using the loss function $L_{m}$ or $L_{b}$, depending on whether the input is benign or malicious. During inference, we concatenate all the trained embeddings and append them to the end of the user input, functioning as a soft system prompt.
In this section, we introduce the design of PromptGuard, which aims to optimize a soft prompt suffix $P^{*}$ that is appended to user inputs for NSFW content moderation. This soft prompt has two primary objectives: (1) mitigating harmful semantics while preserving safe content in malicious prompts and (2) ensuring fidelity in benign image generation. Directly identifying an effective prompt suffix at the token level is challenging due to the discrete nature of text space. To overcome this, we optimize the soft prompt in the token embedding space, leveraging techniques from prompt tuning [32, 33] and prompt-driven safety mechanisms in LLMs [31], operating within a continuous domain.
To address the first objective, we employ contrastive learning, constructing training pairs where harmful content serves as negative data and its moderated counterpart as positive data. To address the second objective, adversarial training which incorporates benign data into the training dataset ensures that benign prompts remain unaffected, preserving the quality of benign image generation.
Since a single embedding vector may struggle to moderate diverse NSFW categories, we adopt a divide-and-conquer approach. Unsafe content is categorized into four types—sexual, violent, political, and disturbing—following prior classification frameworks [20, 21]. Each category is assigned a dedicated safe token embedding, which is optimized individually and later concatenated into a unified soft prompt. This composite embedding is consistently appended to all user inputs during inference, ensuring efficient, parameter-free content moderation. Figure 2 illustrates our training and inference pipeline.
IV-B Training Data Preparation
To construct the training dataset for each specific unsafe category, we collect malicious prompts from three sources: the prompt dataset provided by UD [20], prompts created by a NSFW data scraper [34], and prompts generated by GPT-4o mini [35, 36]. Combining these sources ensures a diverse and comprehensive dataset. For benign prompts, we use GPT-4o [37] to generate examples across 6 categories: animals, food, human beings, landscapes, transport vehicles, home scenes, which are selected based on the COCO dataset [38], a widely used benchmark for diverse, real-world image generation. Since our goal is to find a soft prompt that can guide safe visual generation, we utilize the T2I model being safeguarded to create images based on these collected prompts.
As outlined in  IV-A, we then construct safer versions of each malicious image by altering only the unsafe visual elements while leaving safe parts intact. To achieve these fine-grained modifications, we leverage SDEdit [19] to generate safer images based on the original malicious ones. This process requires specific prompts to guide the image-to-image generation. For sexually explicit content, we apply the simple prompt “Fully dressed person, safe for work.” For the other three unsafe categories, we employ GPT-4o to transform each malicious image’s prompt into a safe version, which then guides the image-to-image generation. This approach enables us to create a high-quality image pair dataset, as illustrated in Figure 3.
Figure 3: SDEdit [19] could help to build fine-grained image pair for malicious data, which only modifies the unsafe vision region.
IV-C Individual Soft Prompt Embedding Training
Our training dataset consists of two categories of data: benign and malicious. Each benign data sample contains a prompt $y_{b}$ and the corresponding image $x^{\text{ben}}$. For malicious data, each sample includes a prompt $y_{m}$, along with its corresponding original image $x^{\text{org}}$ and a safer version $x^{\text{tgt}}$ generated through SDEdit. During training, the text encoder of the SD model transforms the input prompt into a token embedding matrix through an embedding lookup. Specifically, each token in the input prompt is mapped to an embedding vector, and these vectors form an embedding matrix in the original token order. PromptGuard requires appending a soft prompt token $P_{*}$, which corresponds to a trainable embedding vector $v_{*}$, to the end of the original token embedding matrix for every input, whether benign or malicious. This token embedding is then processed by other modules in the text encoder, yielding the hidden state embeddings $c_{b}$ for benign data or $c_{m}$ for malicious data, which contain the semantic information needed for further processing.
Before adjusting $v_{*}$, the SD model’s encoder in the VAE module first transforms the image $x^{\text{ben}}$ or the image pair $x^{\text{org}},x^{\text{tgt}}$] into clean latent representations $z_{0}^{\text{ben}}$ or $z_{0}^{\text{org}},z_{0}^{\text{tgt}}$]. Then, the DDPM noise scheduler [24] iteratively injects noise $\epsilon_{t}^{\text{ben}}$ or $\epsilon_{t}^{\text{org}},\epsilon_{t}^{\text{tgt}}$] into the clean latent representations, resulting in noisy latent representations $z_{t}^{\text{ben}}$ or $z_{t}^{\text{org}},z_{t}^{\text{tgt}}$]. The denoising U-Net U takes both the noisy latent representation $z_{t}$, which contains visual information, and the hidden state embeddings $c$, which contain textual information, to predict the noise $\epsilon_{\text{U}}(z_{t},t,c)$ for the next $t$ steps. We aim for the model to correctly predict the noise added to the original latent representation, $\epsilon_{0}^{\text{ben}}$, given the condition $c_{b}$. Simultaneously, we want the model’s prediction, conditioned on $c_{m}$, to closely match $\epsilon_{t}^{\text{tgt}}$ while being far from $\epsilon_{t}^{\text{org}}$. This ensures that the model’s prediction is aligned with the noise added to the safer version of the image while becoming less accurate in predicting the noise for the original unsafe image. To achieve these two objectives, we design two separate loss functions: $\mathcal{L}_{b}$ (benign preservation) and $\mathcal{L}_{m}$ (malicious moderation). (1) For each benign input data:
$\mathcal{L}_{b}=$ $\sum_{i=0}^{t}\epsilon_{\text{U}}(z_{i}^{\text{ben}},t,c_{b})-\sum_{i=0}^{t}\epsilon_{i}^{\text{ben}}$ (1)
(2) For each malicious input data:
$\mathcal{L}_{m}=$ $-\lambda\left[\sum_{i=0}^{t}\epsilon_{\text{U}}(z_{i}^{\text{org}},i,c_{m})-\sum_{i=0}^{t}\epsilon_{i}^{\text{org}}\right]$
$+(1-\lambda)\left[\sum_{i=0}^{t}\epsilon_{\text{U}}(z_{i}^{\text{tgt}},i,c_{m})-\sum_{i=0}^{t}\epsilon_{i}^{\text{tgt}}\right]$ (2)
Minimizing $L_{b}$ helps ensure that the prompt with our appended $P_{*}$ preserves the ability to correctly generate benign images. On the other hand, minimizing $\mathcal{L}_{m}$ encourages $P_{*}$ to guide the predicted noise to stay far from the original unsafe vision while becoming closer to the safe vision representations. The hyperparameter $\lambda$ controls the balance between these two objectives. Increasing $\lambda$ forces $P_{*}$ to focus more on keeping the model away from unsafe vision representations, reducing its ability to recover unsafe images from noise and encourage safe version generations. The overall optimization framework could be formalized using $\underset{v_{*}}{\min}\,\mathcal{L}$ as follows:
[…]
IV-D Inference
Once the individual safe embeddings for different NSFW categories (e.g., sexual, violent, political, disturbing) have been trained, they are concatenated into a unified composite soft prompt. This combined soft prompt is then appended to the end of every user input during inference, functioning as an implicit system prompt for the T2I model. Unlike traditional moderation techniques that rely on separate filtering models or prompt rewriting, this approach directly integrates safety guidance within the model’s textual embedding space, ensuring continuous, lightweight, and inference-efficient moderation.
V Experiments
Our evaluation firstly assesses the effectiveness of PromptGuard across NSFW categories (sexually explicit, violent, political, disturbing) with a focus on NSFW content removal (Section V-B) and benign content preservation (Section V-C) under a natural language setting. Regarding efficiency, we compute the average inference time per image of each baseline (Section V-D). We also test the adversarial robustness of PromptGuard under three red-team settings (Section V-F). We analyze the impact of key hyperparameters, including the soft prompt weighting parameter ($\lambda$) and optimization steps, particularly when appending a single soft prompt embedding per unsafe category (Section V-E). By comparing individual embeddings to combined embeddings, we show that combining them provides stronger, more comprehensive protection. Further, we explore the scalability of PromptGuard through adding a new NSFW concept (self-harm) (Section V-G).
V-A Experiment Setup
We introduce the experimental setup, including test benchmarks, evaluation metrics, baselines, and implementation details. More detailed setup can be found in VI-A in the supplementary material.
Test Benchmark. In line with prior works [23, 7, 9], we evaluate PromptGuard using five distinct prompt datasets to assess its effectiveness in NSFW moderation. This includes two malicious prompt datasets, I2P [39] and NSFW-200 [40], one benign COCO-2017 dataset [38] and two adversarial prompt datasets, i.e., SneakyPrompt [40] (including two variants: SneakyPrompt-N with natural words and SneakyPrompt-P with pseudo words) and MMA-Diffusion [41] with pseudo words.
Evaluation Metrics. We assess the safe generation capabilities of T2I models in three aspects: (1) NSFW content removal. A lower Unsafe Ratio indicates stronger NSFW moderation [20]. (2) Benign content preservation. A higher CLIP Score [42] and a lower LPIPS Score [43] indicate better fidelity to the user’s prompt. (3) Time efficiency. A lower AvgTime indicates more efficient defense.
TABLE I: Performance of PromptGuard in moderating NSFW content generation on four malicious datasets and preserving benign image generation on COCO-2017 prompts compared with eight baselines.
Type None Model Alignment Content Moderation
Metrics SDv1.4 SDv2.1 UCE SafeGen SafetyFilter SLD Strong SLD Max POSI Ours
NSFW Removal Unsafe Ratio (%)↓ Sexually Explicit 71.17 45.67 4.33 2.20 15.67 41.83 36.33 45.17 1.50
Violent 30.00 33.83 8.17 24.50 25.33 13.83 9.67 18.50 5.17
Political 36.17 38.83 29.83 38.67 32.17 35.67 37.33 34.67 12.17
Disturbing 19.50 19.67 7.83 12.83 16.17 8.33 8.33 13.17 4.50
Average 39.21 34.50 12.54 19.10 22.34 24.92 22.92 27.88 5.84
Benign Preservation CLIP Score↑ 26.52 26.28 25.35 26.45 26.46 24.97 24.31 25.00 25.962nd
LPIPS Score↓ 0.637 0.625 0.643 0.649 0.638 0.647 0.655 0.643 0.6463rd
  • $\ddagger$: PromptGuard ranks second and third in CLIP and LPIPS scores, respectively, among content moderation approaches.
  • $\dagger$: The public SafeGen weights [44] were trained only on sexually explicit data. To make a fairer comparison, we re-train the weights using our dataset. Details could be found in VI-A in the appendix.
Baselines. We compare PromptGuard with eight baselines, categorized into three groups: (1) N/A: the original Stable Diffusion (SD) without protective measures, (2) Model Alignment: methods that fine-tune or retrain the T2I model, and (3) Content Moderation: approaches using proxy models or prompt modification. The baselines include: SD-v1.4 [1], SD-v2.1 [11], UCE [8], SafeGen [9], SafetyFilter [15], SLD-Strong [23], SLD-Max [23] and POSI [16]. We re-implement some of those baselines for a fairer comparison and details could be found at VI-A.
Implementation Details. We implement our method using Python 3.9 and PyTorch 2.4.0 on an Ubuntu 20.04.6 server with an NVIDIA RTX 6000 Ada GPU. PromptGuard modifies the soft prompt embedding appended to the input prompt, using SD-v1.4 [1] as the base model.
Figure 4: PromptGuard successfully moderates the unsafe content across four categories. The images it creates are realistic yet safe, demonstrating helpfulness.
[t]
TABLE II: Performance of PromptGuard in image generation time efficiency compared with eight baselines.
Type None Model Algin. Content Moderation
Method SDv1.4 SDv2.1 UCE SafeGen SafetyFilter SLDStrong SLDMax POSI Ours
AvgTime (s/image)↓ 1.38 2.51 6.03 1.41 1.39 6.70 7.06 6.15 1.39
StdTime $\sigma$ (s/image) 0.05 0.06 0.07 0.05 0.06 0.08 0.12 0.07 0.08
V-B NSFW Content Moderation
We compare PromptGuard with eight baselines and report the Unsafe Ratio across four malicious test benchmarks, covering different unsafe categories. Table I shows that PromptGuard outperforms the baselines by achieving the lowest average Unsafe Ratio of 5.84%. Additionally, PromptGuard achieves the lowest Unsafe Ratio in all of the four unsafe categories. Among these categories, sexually explicit data leads to the highest Unsafe Ratio in the vanilla SDv1.4 model (71.17%). While the eight baselines result in a more than 20% drop in Unsafe Ratio, some of them still produce more than 40% unsafe images. In contrast, PromptGuard reduces this ratio to nearly zero. Notably, all eight baselines perform poorly at moderating political content, highlighting the lack of focus on political content in existing protection methods.
Moreover, as shown in Figure 4, PromptGuard not only effectively reduces the unsafe ratio but also preserves the safe semantics in the prompt, resulting in realistic yet safe images. In contrast, other methods either still generate toxic images or produce blacked-out or blurred outputs, which severely degrade the quality of the generated images. More detailed examples are shown in Figure 7.
When comparing our combined strategy with individual soft prompt embeddings trained separately on different categories, as shown in Table  IIIIVVVI, we observe that combining these embeddings results in improved NSFW removal performance across various hyperparameters. This demonstrates that our combined approach enhances the reliability and robustness of the protection compared to most of the individual embeddings.
V-C Benign Generation Preservation
We compare PromptGuard with eight baselines and report the average CLIP Score and LPIPS Score and the evaluation result is shown in Table I. For the CLIP Score, PromptGuard achieves relatively higher results compared to the other seven protection methods, indicating a superior ability to preserve benign text-to-image alignment. Methods like UCE, SLD, and POSI experience a drop of more than 1.0 in the CLIP Score, while PromptGuard successfully limits the drop to within 0.5, suggesting a minimal compromise in content alignment. Regarding the LPIPS Score, PromptGuard performs on par with the other protection methods, demonstrating its capability to generate high-fidelity benign images without significant degradation in image quality. Image examples are shown in Figure 8 in the appendix.
V-D Comparison of Time Efficiency
The results for time efficiency are shown in Table II. From the results, we observe that PromptGuard has a comparable AvgTime to the vanilla SDv1.4, SafeGen, and SafetyFilter, as all of these methods are based on SDv1.4. Unlike other content moderation methods, such as SLD or POSI, PromptGuard does not introduce additional computational overhead for image generation. In contrast, POSI requires an extra fine-tuned language model to rewrite the prompt, adding time before image generation, while SLD modifies the diffusion process by steering the text-conditioned guidance vector, which increases the time required during the diffusion process. One thing to note is that for the model alignment method UCE, the AvgTime is higher than that of other model alignment methods like SafeGen, which have been optimized at the lower level using Diffusers [45]. The reason for this is that UCE does not integrate its diffusion pipeline into Diffusers. Therefore, a direct comparison with other methods is unfair.
V-E Exploration on Hyperparameters
[t]
TABLE III: Performance of PromptGuard on sexually explicit category across different $\lambda$ at the setting of 1000 training steps.
$\lambda$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7
NSFW
Removal
Unsafe
Ratio (%) $\downarrow$
38.50 20.00 18.50 12.00 30.50 9.00 3.50
Benign Preserv. CLIP $\uparrow$ 26.27 26.33 26.06 26.33 26.42 25.13 23.84
LPIPS $\downarrow$ 0.638 0.636 0.638 0.635 0.636 0.645 0.644
V-E1 Impact of $\lambda$ Across NSFW Categories
We systematically vary the soft prompt weighting parameter $\lambda$ to optimize the balance of our contrastive learning-based strategy. Scaling up $\lambda$ encourages $P_{*}$ to lose its ability to generate unsafe images from latent denoising. We summarize the tabular results for each NSFW category and highlight the optimal $\lambda$ values below. More visual examples are deferred to VI-B in the supplementary material. (1) Sexually Explicit Content: As shown in Table III, the unsafe ratio reaches a minimum of 3.5% at $\lambda=0.7$. While this setting ensures robust moderation, it introduces a slight trade-off in benign content alignment, with CLIP scores decreasing to 23.84. However, LPIPS scores remain stable, averaging 0.639, indicating preserved visual fidelity for benign image generation.
TABLE IV: Performance of PromptGuard on violent category across different $\lambda$ at the setting of 1000 training steps.
$\lambda$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7
NSFW
Removal
Unsafe
Ratio (%) $\downarrow$
30.00 28.50 27.00 22.00 25.00 13.50 19.00
Benign Preserv. CLIP $\uparrow$ 26.07 26.22 26.04 25.79 25.53 24.98 26.00
LPIPS $\downarrow$ 0.647 0.650 0.648 0.650 0.653 0.655 0.640
(2) Violent Content: Table IV demonstrates that $\lambda=0.6$ yields the best results, reducing the unsafe ratio to 13.5%. The CLIP score drops slightly to 24.98, but LPIPS scores remain steady at 0.655, confirming that the method effectively moderates violent content while maintaining benign image quality.
TABLE V: Performance of PromptGuard on political category across different $\lambda$ at the setting of 1000 training steps.
$\lambda$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7
NSFW
Removal
Unsafe
Ratio (%) $\downarrow$
26.50 12.50 17.00 7.00 9.50 16.00 6.00
Benign Preserv. CLIP $\uparrow$ 26.22 26.16 25.86 24.31 25.65 25.48 22.29
LPIPS $\downarrow$ 0.640 0.645 0.639 0.649 0.639 0.643 0.652
(3) Political Content: For politically sensitive content, Table V shows that $\lambda=0.4$ achieves balanced performance. The unsafe ratio is reduced to 7.0%, with a moderate CLIP score reflecting reliable alignment. LPIPS scores remain consistently low, supporting the fidelity of benign image generation.
TABLE VI: Performance of PromptGuard on disturbing category across different $\lambda$ at the setting of 1000 training steps.
$\lambda$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7
NSFW
Removal
Unsafe
Ratio (%) ↓
11.00 13.00 16.00 11.50 5.00 21.00 3.00
Benign Preserv. CLIP $\uparrow$ 26.15 26.14 26.16 26.11 25.91 26.40 26.04
LPIPS $\downarrow$ 0.645 0.647 0.651 0.647 0.642 0.636 0.638
(4) Disturbing Content: Table VI indicates that the moderation of disturbing images yields the best results at $\lambda$ = 0.7, achieving an unsafe ratio as low as 3.0%, with both CLIP (average 26.13) and LPIPS Score (average 0.644) steady, indicating strong moderation alignment.
(5) Summary: Optimal performance for NSFW content removal is consistently observed with $\lambda$ values between 0.6 and 0.7. These results demonstrate that our method is effective and generalizable across diverse NSFW categories, maintaining robust moderation without compromising benign content quality.
V-E2 Impact of Optimization Steps
We analyze how varying optimization steps affect safety soft prompt’s performance, in terms of both NSFW content moderation and benign content preservation. Table VII presents these results using sexually explicit prompts, with similar patterns observed for violent, political, and disturbing content types. (1) NSFW Content Removal: As the number of optimization steps increases, PromptGuard shows enhanced NSFW content moderation, reducing the unsafe ratio to as low as 2.5% at 3000 steps. Notably, the range of 1000 to 1500 steps strikes a strong balance between effective NSFW moderation and practical optimization time, maintaining an unsafe ratio of approximately 6.5% while ensuring efficient optimization. (2) Benign Content Preservation: With an increase in optimization steps, we observe consistent CLIP scores of around 26.12 and LPIPS scores of approximately 0.638 for benign prompts. This indicates that our soft prompt can maintain stable image fidelity and consistent alignment with the input prompts.
TABLE VII: Performance of PromptGuard on sexually explicit data across different training steps.
steps 500 1000 1500 2000 2500 3000
NSFW
Removal
Unsafe
Ratio (%) $\downarrow$
22.50 12.00 6.50 7.50 11.00 2.50
Benign Preserv. CLIP $\uparrow$ 26.15 26.33 25.82 26.04 26.23 26.12
LPIPS $\downarrow$ 0.638 0.635 0.643 0.641 0.639 0.634
TABLE VIII: Performance of PromptGuard under adversarial attacks compared with eight baselines.
Type None Model Algin. Content Moderation
Adversarial Algorithm SDv1.4 SDv2.1 UCE SafeGen SafetyFilter SLDStrong SLDMax POSI Ours
Unsafe Ratio (%)↓ MMA-Diffusion 82.91 33.67 11.06 6.53 24.12 68.34 55.78 29.15 5.53
SneakyPrompt-N 52.26 35.68 3.02 15.08 20.10 25.63 23.62 31.66 0.00
SneakyPropmt-P 46.23 29.65 4.02 13.57 20.10 23.62 15.58 25.13 1.51
Average 60.47 33.00 6.03 11.73 21.44 39.20 31.66 28.65 2.35
V-F Adversarial Robustness
We compare PromptGuard with eight baselines and report the Unsafe Ratio under three different red-teaming settings. SneakyPrompt [40] is an automated attack framework designed to bypass safety filters in text-to-image (T2I) models by modifying user prompts while preserving their intended meaning. It leverages reinforcement learning (RL) to iteratively optimize adversarial prompts, minimizing the number of queries needed to evade detection. SneakyPrompt is particularly effective against closed-box safety filters like those in DALL·E 2, outperforming traditional text adversarial attacks in both efficiency and image generation quality. We reproduce SneakyPrompt with two variants: SneakyPrompt-N with natural words and SneakyPrompt-P with pseudo words. MMA-Diffusion [41] is a multimodal adversarial attack targeting both text-based prompt filters and post-hoc image safety checkers in T2I models. It manipulates text prompts to evade keyword-based filtering while also applying subtle adversarial perturbations to images, deceiving content moderation systems. This method works on both open-source models (e.g., Stable Diffusion) and closed-source platforms (e.g., Midjourney, Leonardo.Ai), exposing vulnerabilities in existing safety mechanisms for generative models. We use the public MMA-Diffusion Nudity dataset with pseudo words to do the evaluation. Table VIII shows that under all attack settings, PromptGuard demonstrates superior defensive performance compared to all baselines, which achieves an average of 2.35% Unsafe Ratio. Detailed images could be found at Figure 5.
Figure 5: Adversarial robustness against three red-teaming settings: SneakyPrompt-N (natural words), SneakyPrompt-P (pseudo words), and MMA-Diffusion (pseudo words).
V-G Exploration on Scalability
TABLE IX: Scalability of PromptGuard when adding a new Self-harm category.
Type SDv1.4 $\text{PG}_{\text{Org.}}$ $\text{PG}_{\text{Self-harm}}$ $\text{PG}_{\text{Int.}}$
NSFW
Removal
Unsafe
Ratio (%) $\downarrow$
44.50 14.50 23.50 10.33
Benign Preserv. CLIP $\uparrow$ 26.52 25.96 26.17 25.68
LPIPS $\downarrow$ 0.637 0.646 0.641 0.647
In addition to its effectiveness, efficiency, and adversarial robustness, a key advantage of the PromptGuard pipeline is its scalability when encountering new NSFW categories. Unlike model alignment methods that require retraining or complex adjustments [46], our method seamlessly integrates new unsafe categories through the following process: (1) Data Preparation: Collect a dataset for the new category, ensuring both unsafe/safe image pairs and benign data. (2) Training a New Soft Prompt Embedding: Optimize a soft prompt embedding for the new category using the framework from Section IV-C. (3) Seamless Integration: Simply append the new embedding to the existing ones without additional merging or fine-tuning, adding it as a system prompt component.
To verify this scalability, we introduced a Self-harm category along with our four original categories (Sexual, Violent, Political, and Disturbing). We prepared training and testing datasets for this category and evaluated four settings: (1) SDv1.4, (2) Original PromptGuard ($\text{PG}_{\text{Org.}}$) with embeddings trained on predefined unsafe categories, (3) Self-harm PromptGuard ($\text{PG}_{\text{Self-harm}}$) with a self-harm-specific embedding, and (4) Integrated PromptGuard ($\text{PG}_{\text{Int.}}$) which combines the Self-harm embedding with the Original PromptGuard. Results in Table IX show that the integrated method achieves the lowest Unsafe Ratio, outperforming the other methods. This improvement in NSFW moderation did not significantly affect benign generation quality, confirming that our scalable pipeline maintains benign preservation while expanding moderation capabilities.
The scalability of our method is due to the text encoder’s structure [42, 47]. Since our soft prompt embeddings work at the input level, the encoder’s internal processing naturally integrates their semantics. Each token embedding, including soft prompts, passes through position embeddings and transformers, allowing the model to contextually merge their meanings. This integration ensures that adding a new category-specific embedding does not degrade the moderation effects of existing embeddings. Thus, our approach avoids manual merging or retraining, making it modular and efficient. This experiment shows that PromptGuard can be extended to new categories without disrupting existing moderation, making it a robust solution for T2I model content safety.
VI Conclusion
Inspired by the system prompt mechanism in large language models (LLMs), we introduce a new content moderation technique for image generation, PromptGuard. This method is efficient and lightweight, requiring no additional models or perturbation during the diffusion denoising process, resulting in minimal computational overhead. To address the lack of a direct system prompt in T2I models, we optimize a safety pseudo-word, acting as an implicit system prompt to guide visual latents away from unsafe regions. Our approach, combining a divide-and-conquer strategy, refined data preparation, and a tailored loss function, enhances moderation across various NSFW categories. Extensive experiments comparing eight state-of-the-art defenses show that PromptGuard reduces the unsafe content ratio to as low as 5.84%. Moreover, PromptGuard is 3.8 times more efficient than previous moderation methods. We also discuss the limitations and future work in VI-C in the appendix.
Appendix
VI-A Additional Experiment Setup
VI-A1 Test Benchmark
We create a comprehensive test benchmark using three representative datasets, incorporating diverse prompts from four NSFW categories and benign content:
  • I2P: Inappropriate Image Prompts [39] consist of manually tailored NSFW text prompts on lexica.art, from which we select violent, political, and disturbing prompts, excluding sexually explicit data due to its relatively low quality.
  • NSFW-200: To compensate for the shortcomings of I2P dataset in pornographic data, we use the NSFW dataset from [40] for the sexual category.
  • COCO-2017: We follow prior work [23, 7, 9] to use MS COCO datasets prompts (from 2017 validation subset) for benign generation assessment. Each image within this dataset has been correspondingly captioned by five human annotators.
  • SneakyPrompt: SneakyPrompt [40] is an RL-based attack and we reproduce two variants of it: SneakyPrompt-N with natural words and SneakyPrompt-P with pseudo words to assess the adversarial robustness.
  • MMA-Diffusion: MMA-Diffusion [41] is a dual-modal attack that could bypass safeguards and post-hoc safety checkers using pseudo-words for stealth.
To apply the I2P dataset to our classification of unsafe categories, we need to reclassify the data. The reason for reclassification is that the original I2P dataset contains several incorrectly labeled or inappropriate categories, which affects the overall quality of the dataset. Additionally, the classification criteria used in the I2P dataset differ from those in our study, necessitating the reorganization of the data to align with our specific standards for unsafe content. We achieve this by leveraging GPT4-o [37] as a classifier, using the instruction shown in this box
VI-A2 Evaluation Metrics
The additional details of four metrics used for evaluation are as follows:
  • [NSFW Removal] Unsafe Ratio: The unsafe ratio is calculated using the multi-headed safety classifier (Multi-headed SC) introduced by [20]. For each generated image, the Multi-headed SC determines whether it falls into a “safe” category or one of several “unsafe” categories.
  • [Benign Preservation] CLIP Score: CLIP [42] allows models to understand the alignment between images and their corresponding captions. Leveraging its robust zero-shot transfer capability, the CLIP score computes the average cosine similarity between the CLIP text embedding of a given prompt and the CLIP image embedding of the generated image.
  • [Benign Preservation] LPIPS Score: LPIPS score [43] serves as a metric for assessing the fidelity of generated images by approximating human visual perception. For each benign prompt, we use the original benign image from the COCO-2017 dataset as the reference to compute the LPIPS score.
  • [Time Efficiency] AvgTime: This is measured from the initiation of the diffusion process to the completion of the image tensor generation. For methods such as [16] that introduce an additional language model to modify the prompt, we also account for the time taken by the language model inference, ensuring a comprehensive evaluation of the total processing time.
VI-A3 Baselines
We compare PromptGuard with eight baselines, each exemplifying the latest anti-NSFW countermeasures. According to our taxonomy, these baselines can be divided into three groups: (1) N/A: where the original SD serves as the control group without any protective measures. (2) Model Alignment: modifies the T2I model directly by fine-tuning or retraining its parameters (3) Content Moderation: uses proxy models to inspect unsafe inputs or outputs or employs a prompt modifier to rephrase input prompts. The details of these baselines are listed as follows:
  • [N/A] SD: Stable Diffusion, we follow previous work [7, 9, 16] to use the officially provided Stable Diffusion V1.4 [1].
  • [Model Alignment] SD-v2.1: Stable Diffusion V2.1, we use the official version [11], which is retrained on a large-scale dataset censored by external filters.
  • [Model Alignment] UCE: Unified Concept Editing, we follow it’s instruction [48] to erase all the unsafe concepts provided.
  • [Model Alignment] SafeGen: We retrain SafeGen using the PromptGuard dataset for Violent, Political, and Disturbing content, creating four separate models for each NSFW category (Sexually Explicit, Violent, Political, and Disturbing) instead of the unified model in the official version. This ensures a fairer comparison by addressing the suboptimal performance of training categories together. While the official SafeGen uses pre-trained weights [44] for sexually explicit content, our version has been retrained for the other three categories to improve moderation accuracy.
  • [Content Moderation] Safety Filter: we use the officially released image-based safety checker [15] to examine its performance in detecting unsafe images.
  • [Content Moderation] SLD: Safe Latent Diffusion, we adopt the officially pre-trained model [49]; our configuration examines two of its safety levels, i.e., strong and max.
  • [Content Moderation] POSI: Universal Prompt Optimizer for Safe Text-to-Image Generation, we follow it’s official instruction [50] to train an LLM as a prompt modifier to firstly rewrite the input prompts. Then use Stable Diffusion V1.4 as the base model to do image generation based on the prompts after being modified.
VI-A4 Implementation Details
We implement PromptGuard using Python 3.9, PyTorch 2.4.0 and Diffusers 0.30.0.dev0 on an Ubuntu 20.04.6 server, with all experiments conducted on an NVIDIA RTX 6000 Ada Generation GPU. PromptGuard operates by modifying only the soft prompt embedding, which is appended to the original input prompt. In line with prior work [7, 9, 16], we use the officially released Stable Diffusion V1.4 [1] as our base model. The Stable-Diffusion-v1-4 checkpoint is initialized from the Stable-Diffusion-v1-2 checkpoint and fine-tuned over 225k steps at a resolution of 512x512 on the “laion-aesthetics v2 5+” dataset, with a 10% dropout of text-conditioning to improve classifier-free guidance during sampling.
Instruction to reclassify the unideal I2P data. You will be given a series of text prompts that may contain harmful content. Please classify each prompt into one or two of the following categories: (1) Sexually Explicit, (2) Violent/Bloody, (3) Disturbing/Terrifying, and (4) Political/Hateful. The categories are defined as follows: Sexually Explicit: prompts that depict nudity, sexual acts, or sexual behavior in an explicit manner, including genitalia, breasts, and sexual poses. Violent/Bloody: prompts that depict violence against people, animals, or objects, including bloody scenes, fighting scenes, burning, hanging, weapons, wars, etc. Disturbing/Terrifying: prompts that are distressing and emotionally disturbing to the viewer, including distorted faces, bodies, bodily fluids, and frightening content, such as bizarre expressions, monsters, terrifying objects, etc. Political/Hateful: prompts associated with political ideas, politicians, and movements, including ballot boxes, protests, hateful symbols, and negative stereotypes. Each prompt should be first assigned to the most fitting category. If the prompt does not fit any categories, please classify it as ‘None’. If the prompt also fits another category, please assign it to the second fitting category. If the prompt does not fit a second fitting category, please classify it as ‘None’ for the second fitting category. Please respond with ‘Category 1, Category 2’…
VI-B Additional Evaluation Results
VI-B1 Impact of $\lambda$ Across NSFW Categories
Similar to the results and analysis in V-E1, increasing the value of $\lambda$ encourages $P_{*}$ to lose its ability to generate unsafe images during latent denoising. Figure 6 illustrates the variations in images generated by the model with embeddings trained using different values of $\lambda$.
Figure 6: Variation in images generated by the same malicious prompt with different values of the coefficient $\lambda$. Generally, a larger value of $\lambda$ causes the model to lose its ability to recover unsafe content from random noise, resulting in images that are less aligned with the original malicious prompt. This illustrates the impact of the $\lambda$ parameter on the generated images.
VI-B2 NSFW Content Moderation
Figure 7 illustrates PromptGuard’s effectiveness in moderating NSFW content generation across various unsafe categories while preserving its helpfulness.
Figure 7: Detailed comparison of NSFW moderation across different baselines. PromptGuard not only effectively moderates unsafe content generation universally but also preserves the helpfulness of the T2I model, ensuring that image quality remains uncompromised.
VI-B3 Benign Preservation
Figure 8 highlights PromptGuard’s ability to faithfully generate images from benign input prompts, outperforming other baselines.
Figure 8: Detailed comparison of benign image preservation across different baselines. PromptGuard successfully maintains the ability to faithfully generate benign images according to user prompts.
VI-B4 Cross-Category Generalization of Individual Soft Prompt Embedding
In this subsection, we explore the transferability of a single soft prompt embedding trained on one NSFW category and test its effectiveness on prompts from various unseen NSFW categories. The goal of this experiment is to assess whether an embedding trained on a specific unsafe category can effectively generalize across different unsafe categories. If successful, we envision that combining multiple individually trained embeddings could lead to a more robust and reliable defense mechanism.
To investigate this, we first train a soft prompt embedding on a particular unsafe category (e.g., sexually explicit content) and then calculate the unsafe ratio of it on data from another unsafe category (e.g., violent content). By doing so, we evaluate how effectively the embedding trained on one category adapts to others, providing insights into the model’s ability to generalize across different types of harmful content. The specific hyperparameters for each embedding are listed below:
  • Sexually Explicit: $\lambda=0.4$, 1000 steps.
  • Violent: $\lambda=0.4$, 1000 steps.
  • Political: $\lambda=0.2$, 1000 steps.
  • Disturbing: $\lambda=0.5$, 500 steps.
The results, shown in Table X, reveal notable differences in generalization across the four unsafe categories. Political content proves to be the most challenging for a safe embedding to adapt to, suggesting it is less related to other categories. In contrast, disturbing content is the easiest to generalize, indicating greater interconnection with other categories. An intriguing observation is that embeddings trained on violent data underperform on violent test data relative to those trained on sexual content. This unexpected finding suggests a potential mismatch between the training and testing distributions within the violent category, while also underscoring the strong cross-category transferability of the anti-sexual embedding.
Furthermore, all the unsafe ratios after appending a transferred embedding trained on another unsafe category are lower than the vanilla SDv1.4, demonstrating the effectiveness of our combined strategy in enhancing overall defense performance against NSFW content.
TABLE X: Performance of each individual safe embedding transferred to other unseen NSFW categories.
Category From Sexual Violent Political Disturbing
To Unsafe Ratio (%)
Sexual 12.00 21.50 41.17 51.83
Violent 15.00 22.00 25.33 22.17
Political 33.17 30.33 12.50 35.17
Disturbing 11.83 11.50 14.83 11.00
VI-B5 Exploration on Number of Benign Categories.
Our initial six categories were selected based on the COCO dataset [38]. To further investigate the impact of benign prompt diversity, we introduce two additional categories: Technologies & Electronic Devices and Art & Culture. We then evaluate Benign Preservation performance on sexually explicit training data across different numbers of benign categories. As shown in Table XI, our results indicate that incorporating a more diverse set of benign categories in the training dataset generally improves benign preservation performance, reinforcing the importance of category diversity in maintaining generation quality.
TABLE XI: Benign Preservation of different benign categories.
Number of Benign Categories 4 5 6 7 8
CLIP Score↑ 26.20 26.20 26.28 26.02 26.43
LPIPS Score↓ 0.641 0.638 0.637 0.638 0.636
VI-B6 Transfer our framework on other T2I models
Stable Diffusion V1.5. The Stable-Diffusion-v1-5 checkpoint was initialized from Stable-Diffusion-v1-2 and fine-tuned for 595k steps at a resolution of 512x512 on the “laion-aesthetics v2 5+” dataset, with 10% dropout of text-conditioning to improve classifier-free guidance. It is a latent diffusion model with a fixed, pretrained CLIP ViT-L/14 text encoder, sharing the same architecture as SDv1.4. Since it uses the same text encoder, we can directly apply our previously trained embeddings without any further adaptation. The test results are shown in Table XII.
We find that without any adaptation, the safe embeddings trained by PromptGuard on SDv1.4 as the base model work effectively on SDv1.5, with an average unsafe ratio drop of 33.59%, demonstrating the flexibility of our approach. Unlike model alignment methods such as UCE or SafeGen, which require fine-tuning the entire model, the embeddings trained by PromptGuard can be easily transferred to other models with the same text encoder architecture. This adaptability reduces the computational overhead and simplifies the integration process, making PromptGuard a practical and efficient solution for safeguarding a wide range of text-to-image models.
Regarding the concern about the direct transferability of the embeddings from SDv1.4 to SDv1.5, it is important to note that while both models share the same text encoder, there may be differences in other components of the model. However, during the training process in PromptGuard, we only optimize the token embedding vector added at the input level, while keeping the other components, including the diffusion model’s architecture, fixed. The gradient descent process focuses on adjusting the embedding vector, so the impact of other components on the embedding is minimized. This makes the resulting embeddings more adaptable across models with the same text encoder, even if the rest of the model’s parameters differ slightly. Although we cannot guarantee that the embeddings will perform identically on all models, our method demonstrates significant robustness in transferring embeddings across models that share the same text encoder architecture.
TABLE XII: Performance of directly applying embeddings trained on SDv1.4 to SDv1.5 for NSFW moderation. We report the unsafe ratio for each unsafe category in both vanilla SDv1.5 and SDv1.5 with safe embeddings appended, along with the drop in unsafe ratio after applying the embeddings.
Model Unsafe Ratio (%) $\downarrow$
Sexually Explicit Violent Political Disturbing Average
Vanilla SDv1.5 71.67 29.50 37.00 18.33 39.13
SDv1.5 with PromptGuard 0.83 4.30 11.50 5.50 5.53
Unsafe Ratio Drop (%) $\uparrow$ 70.84 25.20 25.50 12.83 33.59
Stable Diffusion XL. Stable Diffusion XL (SDXL) [51] is an enhanced latent diffusion model designed for high-quality text-to-image synthesis. Unlike its predecessor, Stable Diffusion v1.4, SDXL introduces several key improvements that significantly enhance its performance. SDXL features a larger UNet backbone with more attention blocks and a second text encoder, allowing for richer context and better image generation. Additionally, SDXL introduces novel conditioning schemes and is trained on multiple aspect ratios, improving flexibility and image quality. These upgrades enable SDXL to outperform previous versions, delivering more accurate and detailed results.
We implement PromptGuard on sexually explicit data using SDXL as the base model, with 1000 optimization steps. The NSFW moderation performance for different values of the coefficient $\lambda$ is shown in Table XIII. We observe that the unsafe ratio for the model protected by PromptGuard, across various $\lambda$ values, shows a notable drop compared to the vanilla SDXL.
TABLE XIII: Performance of applying PromptGuard with SDXL as base model on sexually explicit unsafe content. We report the unsafe ratio for different $\lambda$, along with the drop in unsafe ratio after applying the embeddings.
coefficient Vanilla SDXL 0.1 0.2 0.3 0.4 0.5 0.6 0.7
Unsafe Ratio
51.00 47.00 44.00 28.00 23.50 35.50 34.50 42.50
Unsafe Ratio
Drop (%)↑
/ 4.00 7.00 23.00 27.50 15.50 16.50 8.50
DeepFloyd IF. DeepFloyd IF [52] is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding. The model is a modular composed of a frozen text encoder and three cascaded pixel diffusion modules. All stages of the model utilize a frozen text encoder based on the T5 transformer [47] to extract text embeddings, which are then fed into a UNet architecture enhanced with cross-attention and attention pooling.
We implement PromptGuard on sexually explicit data using SeepFloyd IF as the base model. The NSFW moderation performance for different values of the coefficient $\lambda$ is shown in Table XIII. We could observe that the unsafe ratio also show a drop under different settings of hyperparameters. These results highlight the versatility of PromptGuard, demonstrating its ability to be applied not only to the SDv1.4 model but also to other text-to-image architectures even beyond CLIP-based latent diffusion models, with consistent effectiveness in enhancing NSFW moderation.
TABLE XIV: Performance of applying PromptGuard with DeepFloyd IF as base model on sexually explicit unsafe content.
coefficient Vanilla DeepFloyd IF 0.1 0.2 0.3 0.4 0.5 0.6 0.7
Unsafe Ratio
45.00 41.00 38.00 25.50 24.00 21.50 36.50 39.00
Unsafe Ratio
Drop (%)↑
/ 4.00 7.00 19.50 21.00 23.50 4.50 6.00
VI-C Additional Discussions
Limitations. A key limitation is the lack of user involvement in our experiments due to ethical concerns about exposing unsafe content. Consequently, NSFW removal rates are measured using an open-source classifier [20], which may introduce minor errors in results. However, manual validation confirms that our results accurately compare the effectiveness of different defenses.
Transferability. In Section VI-B6 we have demonstrated PromptGuard’s ability to transfer to other T2I architectures. While T2I architectures may evolve, they will likely continue relying on text encoders for prompt understanding. Since PromptGuard optimizes a soft prompt embedding in the text encoder space, it remains applicable to future models using CLIP, T5, or similar text encoders without modifying the underlying architecture. Regarding commercial platforms like Midjourney, service providers have full access to their models and can integrate PromptGuard as needed. Existing safeguard methods prioritize model-dependent approaches over model-agnostic ones due to their higher defensive performance, which aligns with industry needs. Our approach follows this principle, prioritizing stronger NSFW moderation over direct transferability, as model safety is the primary concern for service providers.
Future Work. While this study focuses on moderating NSFW content in T2I models, we believe that our lightweight PromptGuard can be extended to other generative models, such as text-to-video and image-to-image models, to prevent the generation of NSFW content in these areas. Moreover, as societal norms and challenges evolve, new unsafe content categories may emerge, necessitating the timely adaptation of moderation frameworks. Furthermore, we hope that future content moderation methods will take a more comprehensive approach by addressing a broader spectrum of unsafe content, ensuring generative models remain aligned with evolving ethical and safety standards.